OpenStudy (anonymous):

@Mathmale Can you help me please? It takes Brian 15 hours longer to build a model car than it takes John. If they work together, they can build the model car in 4 hours. Using complete sentences, explain each step in figuring out how to determine the time it would take Brian to build the car on his own.

OpenStudy (anonymous):

Its 19 hours. You have upset the mathbreaker with your preference to mathmale. I'd be a disgrace if its wrong. But still ._.

OpenStudy (anonymous):

20 hours sorry *

OpenStudy (anonymous):

How do you know it's 20 hours, though? And I'm sorry to upset you.. .-.

OpenStudy (anonymous):

Hehe just kidding :p Let assume it takes John x hours to complete his car, then it would take brian x+15 hours to complete his. Now if they do it together, they can finish it in 4 hours. So if we look at it differently, it takes john 1/x hours to complete some fraction of the work, and it takes brian 1/(x+15) hours to complete a smaller fraction, and if we add them together, you'll get the work done in one hour, now this must be equal to the work done in 1 hour if they do it together, which is 4 hours. So 1/x + 1/(x+15) = 1/4 now if we solve it, we get x^2 + 15 x = 8x + 60. Now if you solve this, youll get the answer 5, now simply add 15 to it and you get the time taken by brian which is 20 :) all the best

OpenStudy (anonymous):

Oh, thank you, Mathbreaker! :)

OpenStudy (anonymous):

@Mathbreaker how did you get x^2 + 15 x = 8x + 60? That's where I'm confused at

OpenStudy (kaylaprincess):

^^ i dont get that either...